# Inference Optimal VLMs Need Only One Visual Token but Larger Models

### v1

@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have enabled Vision Language Models (VLMs) to perceive, reason, and respond through both text and image inputs (Liu et al., 2023; Alayrac et al., 2022; Dai et al., 2023). Many VLMs are built on top of pretrained vision encoders, like CLIP, and pass the patch-based tokens from the visual encoder into the pretrained LLM backbone at a one-to-one ratio for visual context. This results in the LLM processing hundreds of tokens per image, overshadowing those from the user prompt and accounting for most of inference time compute. Consequently, deploying VLMs in real-world applications, particularly on consumer-side edge devices such as monitoring systems, driving assistants, etc., is often limited by the significant inference cost and resulting latency.
@@
To reduce the inference cost of VLMs, many recent works have focused on decreasing the number of visual tokens, via a small learnable module, prior to passing image tokens into the LLM while minimizing performance degradation (Li et al., 2024c; Shang et al., 2024). For example,  (Li et al., 2024c) learn a cross attention module over the CLIP output tokens to compress the number of tokens. Alternatively, inference FLOPs, proportional to the number of parameters and number of tokens processed, can be reduced by using a smaller LLM. Since both the LLM size and number of visual input tokens directly affect the VLM’s performance, it becomes unclear what the optimal trade-off between the two is. For example, one could process all visual input tokens using a 4B LLM or use an 8B LLM on a reduced set of half the original visual tokens, as both result in similar inference costs — currently, the ideal choice is unknown.
This raises an important question: given a fixed inference budget, what is the optimal trade-off between LLM size and the number of visual tokens processed for downstream performance? In this work, we try to answer this question by building the first inference-time compute-optimal scaling laws for VLMs, modeling performance as a function of both key factors affecting inference cost: LLM size and the number of visual tokens processed. We observe that the downstream error varies 5x faster with LLM parameters than with the number of input visual tokens.
@@
In fact, our scaling laws reveal a striking observation: for visual reasoning tasks, the compute-optimal inference regime entails using the largest feasible LLM with a very small number of visual input tokens — often just one — when the input query can be cached. We show that for any given fixed inference cost, trading off the visual tokens for a larger LLM size leads to a reduction in downstream error for visual reasoning tasks. However, for certain use cases like Optical Character Recognition (OCR) or document understanding tasks, the optimal approach is quite the opposite, requiring as many visual tokens as possible, as token compression proves ineffective for capturing the dense and diverse information present in such tasks.
@@
Most existing work on token compression has focused on reducing visual tokens by a modest factor (e.g., from 576 to 144 tokens or 64 tokens). In contrast, our results underscore the critical importance of pursuing much higher compression rates (e.g., reducing tokens to 1 or 4) for visual reasoning tasks where such compression is not only feasible but also compute-optimal. Building upon these insights, we take initial steps toward developing token compression algorithms specifically tailored for high compression regimes. We propose a query-based token compression approach, recognizing that in extreme compression scenarios, it is essential to selectively curate tokens based on the user’s query to preserve the most relevant tokens. In summary, our work identifies the compute-optimal inference regime for VLMs, emphasizing the importance of high token compression for visual reasoning tasks. We hope these findings will serve as a motivation and foundation for shifting token reduction techniques towards more effective and higher compression ratios.
@@
Our work is organized as follows. We first introduce some preliminaries around inference costs and visual token compression for VLMs in Section 2. Then we talk about our compute optimal scaling laws in Section 3, the results of which motivate our compression algorithm designed for high token-reduction regimes are covered in Section 4.
(a) Scaling laws for VLMs at $Q=0$ (cached text).
(b) Scaling laws for VLMs at $Q=50$ (variable text).
@@
Figure 1: Inference optimal scaling laws for VLMs: The number of visual tokens ($V$) passed to the LLM (after token compression, § 2.2), along with the LLM parameter count ($N$), directly determine the inference cost of VLMs ($\mathcal{O}(N(Q+V))$), where $Q$ is the text input tokens. Since a VLM’s downstream performance is directly affected by both these factors, it is unclear what the optimal trade-off is for a fixed inference compute. In this work, we try to answer this question with our scaling laws. Left (a): We plot the fitted scaling curves, assuming cached text input tokens ($Q=0$). We observe a surprising trend: for visual reasoning tasks, the compute optimal behavior (dotted black curve) requires using a single visual token with the largest possible language model that can fit under the inference budget. Right (b): Inference optimal behavior under $Q=50$ requires slightly higher number of visual tokens as the LLM already incurs a fixed cost due to the text tokens.
@@ 2.1 Estimating Inference cost for VLMs
where $N$ denotes the parameter count of LLM and $T$ denotes the total inference time tokens. We ignore the inference cost stemming from the visual encoder, as we use the CLIP-L vision encoder (Radford et al., 2021) with the same input image resolution across all experiments. In addition, many current open-source VLMs utilize the same encoder, incurring a fixed cost across models.
@@
In the context of VLMs, the total inference tokens, $T$, can be further decomposed as $T=Q+V+G$, where $Q$ represents the text input tokens, i.e., the question/prompt, $V$ represents the number of visual tokens from the vision encoder (after token compression), and $G$ accounts for the generated tokens. In many real world applications, such as driving assistance systems, the text input remains constant (e.g., “Alert the driver if the scene ahead has a hazard”). In these scenarios, the text input can be cached, effectively making $Q=0$ by bypassing self-attention projections and feed-forward calculations. However, in other interactive applications, $Q$ may vary based on dynamic input. We will study the behavior of the downstream error with $FLOPs_{\text{inf}}$ under both $Q=0$ and varying $Q$ regimes. Finally, the generation tokens again are quite small for most inference tasks (single word answers). However, the analysis with increasing $Q$ transfers to increasing $Q+G$ as well.
@@ 2.2 Token Compression in VLMs
As discussed in the previous section, inference FLOPs for VLMs increase proportionally with the number of visual input tokens (e.g., 576 per image with CLIP-ViT-L visual encoder). The number of visual tokens often dominates the total number of tokens processed by the language model, especially in applications where the text input can be cached or is comparatively shorter. Thus, there has been a growing interest in developing approaches to compress the visual information into a fewer number of tokens via a small learnable module. For example, TokenPacker (Li et al., 2024c) learns a small cross-attention module over the image tokens to compress them before passing to the LLM.
@@ 3 Tokens vs Parameters: Inference Time Scaling Laws for VLMs
3.1 Scaling Law Formulation
@@
where $N$ denotes the LLM parameters, $T$ denotes the input visual tokens, $\{A,B,D,\alpha,\beta\}$ are learnable parameters, and $Y(N,T)$ is a measure of model quality. Although traditional scaling laws have been studied in the context of pretraining loss Kaplan et al. (2020), practitioners often use the direct downstream performance to assess model quality (Gadre et al., 2024; Goyal et al., 2024b; Liu et al., 2022). Thus, we use average downstream error on a suite of nine commonly used visual reasoning benchmarks (§ 3.2) as a measure of model quality $Y(N,T)$.
@@
LLM Quality Parameter ($\alpha$): This parameter dictates how the downstream error changes with the complexity of the LLM, i.e., its parameter count. A larger $\alpha$ indicates a better language model, such as Llama3-7B outperforming Llama2-7B, which often stems from better pretraining.
Visual Token Quality Parameter ($\beta$): $\beta$ captures the quality of the visual input tokens fed into the LLM, reflecting the quality of the compression technique. A more effective token compression algorithm would yield a larger $\beta$, allowing for more reductions in number of $T$ visual tokens than less effective methods while maintaining the same downstream performance.
Constants ($A,B,D$): $A\text{ and }B$ are normalizing constants and $D$ refers to irreducible loss, which cannot be reduced even with the largest $N$-sized language model or all $T$ visual tokens (capped at 576 for our choice of vision encoder).
@@ 3.2 Experimental Setup
VLM Training and Evaluation: We use the LLaVA-Next framework (Liu et al., 2024b) to train VLMs with the Qwen-1.5 family of language models as the backbone. Specifically, we utilize the Qwen-$\{0.5,1.8,4,7,14\}$B-chat models (Bai et al., 2023). The pretraining and finetuning dataset and hyperparameters follow Liu et al. (2024a), except we double the effective batch size for finetuning. We use CLIP ViT-L/14 (Radford et al., 2021) as the vision encoder for all experiments, and compress the original 576 tokens to $\{144,64,36,16,4,1\}$ tokens using TokenPacker (Li et al., 2024c).
To estimate the downstream error $Y(N,T)$, we evaluate on 9 commonly used benchmarks for visual reasoning and understanding: MME (Fu et al., 2024), GQA (Hudson & Manning, 2019), AI2D (Kembhavi et al., 2016), MMBench (Liu et al., 2024c), MMMU (Yue et al., 2023), ScienceQA (Lu et al., 2022), MathVista (Lu et al., 2024), POPE (Li et al., 2023c), and ChartQA (Masry et al., 2022). We compute $Y(N,T)$ by averaging the errors of the normalized evaluation metric. For MME, the Cognition and Perception scores were combined and the F1 scores were used for POPE.
Fitting Scaling Laws: We fit the proposed scaling law (Eq. 2) on $\{Y(N,T),N,T\}$ pairs, with $N\in\{0.5B,1.8B,4B,7B\}$ and $T\in\{1,4,16,36,64,144,576\}$ (described in the experiment setup above). We use grid-search, for its stability (Goyal et al., 2024b), to estimate the scaling parameters $\alpha,\beta,A,B,\text{ and }D$. The final scaling law is evaluated on a $N=14B$ VLM model at various number of visual tokens in $T$. Further details about the grid-search fit can be found in Appendix A.2.
@@ 3.3 Results: Estimated Scaling Curves
Recall from Section 2.1 that $FLOPs_{\text{inf}}~{}=\mathcal{O}(N(Q+V))$, where $Q$ represents the input text tokens, and $V$ is the visual input tokens. We first visualize our scaling laws under 2 settings — (a) cached text input (Fig.1(a)): The input text tokens ($Q$) are fixed and can be cached, leading to $FLOPs_{\text{inf}}\sim\mathcal{O}(NV)$, and (b) non-cached text input (Fig.1(b)): The input text tokens are approximated as 50, i.e., $FLOPs_{\text{inf}}=\mathcal{O}(N(50+V))$ (we consider more granular variation of $Q$ in § 3.3.2).
Figure 1 visualizes the fitted scaling curve, illustrating the variation in the average downstream error as inference FLOPs are varied (under both the cached and non-cached text input setting). We vary the inference FLOPs on the x-axis by increasing the number of visual input tokens processed by the LLM (the scatter size), while the color scale indicates the varying number of language model parameters. We make some key observations below.
@@
Consider the change in performance for the 7B model as the number of visual input tokens varies (maroon curves in Fig. 1.) Recent works on visual token compression (Li et al., 2024c; Shang et al., 2024) claim little to no performance degradation with token compression. For example, they report similar performance to the base model’s 576 tokens even when visual token count is reduced to 36 or 144 on certain tasks. However, our scaling curves in Figure 1(a) reveal a different trend, showing a log-linear decrease in visual reasoning performance as the number of visual input tokens is reduced. We believe this discrepancy arises because of the limited downstream evaluation benchmarks considered in the previous works which may not fully capture the VLM’s overall capabilities.
@@
Recall from the scaling law (Eq. 2) that $\alpha$ represents the LLM quality parameter and $\beta$ represents the visual token quality parameter, both denoting the rate at which they influence the downstream error respectively. From Figure 1(a), we observe that $\alpha=0.077$ is more than five times larger than $\beta=0.015$, signifying that VLM error increases significantly faster when reducing the LLM parameters compared to reducing the number of visual tokens. Therefore, when minimizing inference FLOPs, it is more effective to prioritize reducing visual tokens ($V$) first, as the impact on performance is less pronounced than reducing the LLM parameters ($N$). This finding is reflected in Figure 4 where we observe that, under fixed inference compute, using a larger LLM with fewer visual tokens (7B LM w/ 36 tokens) provides better performance than using a smaller LLM with more visual input tokens (1.8B LM w/ 144 tokens) for visual reasoning tasks.
@@
Figure 2: Our scaling laws (fitted on VLMs with 0.5-7B LLMs) estimate the performance of a 14B LLM VLM with an error margin of less than 2%.
@@
We evaluate the accuracy of our scaling laws (fitted on VLMs of 0.5B-7B range) for predicting the performance for larger models. We estimate the performance of Qwen-1.5 14B using our fitted scaling laws. Our scaling laws estimate the performance with an error margin of less than 2%, as visualized in Figure 26(b). The log-linear relationship between the error and number of visual tokens persists, and the greater influence of the LLM’s size compared to visual tokens on performance continues to hold. Thus, for VLMs using 7B language model backbones, it is still optimal to increase LLM size to 14B while reducing visual token count for fixed inference costs.
3.3.1 Compute-Optimal Inference Requires a Single Visual Token
Observe the pareto optimal curve (black dotted curve) in Figure 1(a). For cached query, at any given inference compute, the optimal behavior, i.e., lowest downstream error, occurs when using the largest possible LLM while reducing the number of visual input tokens to one. Thus, for scenarios where the text input can be cached ($Q=0$), such as monitoring systems with static text input, one should utilize the largest LLM possible by reducing the number of visual tokens to fit the inference budget. A similar trend of prioritizing the LLM size holds in the variable text input regime. For example, in Figure 1(b), where the text input length $Q=50$, better performance in a fixed compute budget often results from the larger model with fewer visual tokens with the optimal number of visual tokens is now around 16 (intersection of pareto curve with scaling plot).
This increase in optimal visual tokens as text tokens increase is intuitive, as the VLM incurs a fixed cost for processing the text. Thus, small increases in the number of visual tokens lead to only minor increases in the overall inference cost while improving performance. The key observation is that compute-optimal behavior entails using the largest feasible LLM with very few visual input tokens.
This result has important consequences. Existing literature on token reduction (§ 5) has primarily focused on moderately reducing the number of visual input tokens (e.g., from 576 tokens to 144 or 64 tokens) while trying to match the performance of the base model. However, our results highlight that it is better to operate in a regime with much lower input visual tokens (e.g., 1, 4 or 16), as exchanging visual tokens for larger LLM size reduces the downstream error. This highlights the need to develop token compression techniques tailored for extreme token compression. We take some initial steps in this direction, building on existing token compression algorithms in Section 4.

### v2

# Inference Optimal VLMs Need Fewer Visual Tokens and More Parameters

@@ Abstract
Code available at https://github.com/locuslab/llava-token-compression.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have enabled Vision Language Models (VLMs) to perceive, reason, and respond through both text and image inputs (Liu et al., 2023; Alayrac et al., 2022; Dai et al., 2023). Many VLMs are built on top of pretrained vision encoders, such as CLIP, and pass the patch-based tokens from the visual encoder into the pretrained LLM backbone at a one-to-one ratio for visual context. This results in the LLM processing hundreds of tokens per image, overshadowing those from the user prompt and accounting for most of inference time compute. Consequently, deploying VLMs in real-world applications, particularly on consumer-sided edge devices, e.g., monitoring systems, driving assistants, etc., is often limited by the significant inference cost and resulting latency.
@@
To reduce the inference cost of VLMs, many recent works have focused on decreasing, via merging or pruning, the number of visual tokens passed to the LLM without significant performance degradation (Li et al., 2024c; Shang et al., 2024). Alternatively, inference FLOPs, proportional to the number of parameters and number of tokens processed, can be reduced by using a smaller LLM. Since both the LLM size and number of visual input tokens directly affect the VLM’s performance, it becomes unclear what the optimal trade-off between the two is. For example, a 4B LLM processing all visual input tokens results in similar inference costs to a 8B LLM processing half the number of original visual tokens — currently, the ideal choice is unknown.
@@
This observation raises an important question: given a fixed inference budget, what is the optimal trade-off between LLM size and the number of visual tokens processed for downstream performance? In this work, we answer this question by building the first inference-time compute-optimal scaling laws for VLMs, modeling performance as a function of both key factors affecting inference cost: LLM size and the number of visual tokens processed. Our scaling laws reveal a striking observation: for visual reasoning tasks, the compute-optimal inference regime entails using the largest feasible LLM with a very small number of visual input tokens — usually less than 3% the original number of visual tokens. However, for certain use cases that require detailed image analysis, like Optical Character Recognition (OCR) or document understanding tasks, the optimal approach is quite the opposite, requiring as many visual tokens as possible, as token compression proves ineffective for capturing the dense and diverse information present in such tasks.
@@
Most existing work on token compression focuses on modest reductions (e.g., 576→144 or 64), while our results highlight the importance of much higher rates (e.g., 1–4 tokens) for compute-optimal visual reasoning. Our work identifies the compute-optimal inference regime for VLMs, emphasizing the importance of high token compression for visual reasoning tasks. We hope these findings serve as motivation and foundation for shifting token reduction techniques towards more effective and higher compression ratios.
@@
We organize our work as follows. We first introduce some preliminaries around inference costs and visual token compression for VLMs in Section 2. In Section 3, we formulate and analyze our inference-compute scaling laws, including its generalization across compression techniques and trends across different tasks. We use our inights from these compute optimal scaling laws to motivate and build our compression algorithm designed for high token-reduction regimes in Section 4.
@@
Figure 1: Inference optimal scaling laws for VLMs. The number of visual tokens ($V$) passed to the LLM (after token compression, § 2.2), along with the LLM parameter count ($N$), directly determine the inference cost of VLMs ($\mathcal{O}(N(Q+V))$), where $Q$ is the text input tokens. Since the downstream performance of VLMs is directly affected by both these factors, it makes it unclear what the optimal trade-off is for a fixed inference compute. In this work, we try to answer this question with our scaling laws. Left: We plot the fitted scaling curves, assuming cached text input tokens ($Q=0$). We observe a surprising trend: for visual reasoning tasks, the compute optimal behavior (dotted black curve) requires using a single visual token with the largest possible language model that can fit under the inference budget. Right: Inference optimal behavior under $Q=50$ requires slightly higher number of visual tokens as the LLM already incurs a fixed cost due to the text tokens.
@@ 2.1 Estimating Inference cost for VLMs
where $N$ denotes the parameter count of LLM and $T$ denotes the total inference time tokens. We ignore the inference cost stemming from the visual encoder, as we use the same vision encoder with the same input image resolution across all experiments. In addition, many current open-source VLMs currently utilize the same CLIP-L vision encoder (Radford et al., 2021).
@@
In the context of VLMs, the total inference tokens, $T$, can be further decomposed as $T=Q+V+G$, where $Q$ represents the text input tokens, i.e., the question/prompt, $V$ represents the number of visual tokens from the vision encoder (after token compression), and $G$ accounts for the generated tokens. In many real world applications, such as driving assistance systems, the text input remains constant (e.g., ”Alert the driver if the scene ahead has a hazard”). In these scenarios, the text input can be cached, effectively making $Q=0$ by bypassing self-attention projections and feed-forward calculations. However, in other interactive applications, $Q$ may vary based on dynamic input. We will study the behavior of the downstream error with $FLOPs_{\text{inf}}$ under both the $Q=0$ and varying $Q$ regimes. In our work, we ignore the $G$ term due to our analyzed tasks’ short form responses; however, the analysis with varying $Q$ transfers to $Q+G$ as well.
@@ 2.2 Token Compression in VLMs
As discussed in the previous section, inference FLOPs for VLMs increase proportionally with the number of visual input tokens (e.g., 576 with CLIP-ViT-L visual encoder). Often, the number of visual tokens dominates the total tokens processed by the language model, especially in applications where the text input can be cached or is on the shorter side. Thus, there has been a growing interest in developing approaches to compress the visual information into a fewer number of tokens.
@@ 3 Tokens vs Parameters: Inference Time Scaling Laws for VLMs
3.1 Tokens vs Parameters: Scaling Law Formulation
@@
where $N$ denotes the LLM parameters, $T$ denotes the total inference tokens, $\{A,B,D,\alpha,\beta\}$ are learnable parameters, and $Y(N,T)$ is a measure of model quality. Although traditional scaling laws have been studied in the context of training loss Kaplan et al. (2020), practitioners often use the direct downstream performance to assess model quality (Gadre et al., 2024; Goyal et al., 2024b; Liu et al., 2022). We use average performance error on a suite of nine commonly used visual reasoning benchmarks (§ 3.2) as a measure of model quality $Y(N,T)$.
@@
LLM Quality Parameter ($\alpha$): This parameter dictates how the downstream error changes with the complexity of the LLM, i.e., its parameter count. A higher $\alpha$ indicates a better language model, such as Llama3-7B outperforming Llama2-7B, often due to superior pretraining.
Visual Token Quality Parameter ($\beta$): $\beta$ captures the quality of the visual input tokens fed into the LLM, reflecting the quality of the compression technique. A better token compression algorithm would yield a higher $\beta$, allowing for more reductions of $T$ visual tokens than less effective methods while maintaining the same downstream performance.
Constants $A,B,D$: $A\text{ and }B$ are normalizing constants and $D$ refers to irreducible loss, which cannot be reduced even with the largest $N$-sized language model or all $T$ visual tokens (capped at 576 for our choice of vision encoder).
@@ 3.2 Experimental Setup
VLM Training and Evaluation: We use the LLaVA-Next framework (Liu et al., 2024b) to train VLMs with the Qwen-1.5 family of language models as the backbone. Specifically, we utilize the $\{0.5,1.8,4,7,14\}$B-chat models (Bai et al., 2023). The pretraining and finetuning dataset and hyperparameters follow Liu et al. (2024a), except we doubled the effective batch size for finetuning.
To estimate the downstream error $Y(N,C)$, we test our trained VLMs on a suite of nine commonly used benchmarks for evaluating visual reasoning: MME (Fu et al., 2024), GQA (Hudson & Manning, 2019), AI2D (Kembhavi et al., 2016), MMBench (Liu et al., 2024c), MMMU (Yue et al., 2023), ScienceQA (Lu et al., 2022), MathVista (Lu et al., 2024), POPE (Li et al., 2023c), and ChartQA (Masry et al., 2022). We average the normalized evaluation metric errors to compute $P(N,C)$. For MME, the Cognition and Perception scores were added and normalized, while the F1 score was used for POPE (Liu et al., 2024a). As mentioned in Section 2.1, this set of datasets was selected due to their similar prompt and generation length and overall comprehensiveness.
Visual Token Compression: CLIP ViT-L/14 (Radford et al., 2021) is used as the vision encoder for all experiments, and we compress the original 576 tokens to $\{144,64,36,16,4,1\}$ using one of our variants of TokenPacker (Li et al., 2024c) which replaces interpolation with a convolution for downsampling (no adding query embedding, refer to Section 4 for more details).
Fitting Scaling Laws: We fit the proposed scaling law (Eq. 2) on $\{Y(N,T),N,T\}$ pairs, with $N\in\{0.5,1.8,4,7\}B$ and $T\in\{1,4,16,36,64,144,576\}$. We use grid-search, for its stability (Goyal et al., 2024b), to estimate the scaling parameters $\alpha,\beta,A,B,\text{ and }D$. The final scaling law is evaluated on a $N=14B$ VLM model at various $T$ visual tokens. Further details about the grid-search fit can be found in Appendix A.2.
@@ 3.3 Results: Estimated Scaling Curves
We visualize our scaling laws across inference compute, modeled by $FLOPs_{\text{inf}}~{}=\mathcal{O}(N(Q+V))$ where $Q$ represents the input text tokens, and $V$ is the visual input token, under 2 settings — (a) cached text input (Fig.1 Left) where $Q=0$ and (b) non-cached text input (Fig.1 Right) where $Q=50$. Section 3.3.2 presents a more granular analysis of the effect of $Q$ text input tokens on compute optimal tradeoffs for visual reasoning tasks.
Figure 1 presents the fitted scaling curves, illustrating the variation in average downstream error as a function of inference FLOPs. The scatter sizes represent the number of visual input tokens processed by the language model, while the color scale indicates the varying number of language model parameters. We make some key observations below.
@@
Consider the change in performance for the 7B model as the number of visual input tokens varies (maroon curve in Figure 1.) Recent works on visual token compression (Li et al., 2024c; Shang et al., 2024) claim little to no performance degradation with token compression. For example, they report similar performance to the base model’s 576 tokens even when visual token count is reduced to 36 or 144 on certain tasks. However, our scaling curves in Figure 1 reveal a different trend, showing a log-linear decrease in visual reasoning performance as the number of visual input tokens is reduced. We believe this discrepancy arises because of the limited downstream benchmarks evaluated in previous works which may not fully capture the VLM’s overall capabilities.
@@
Recall from the scaling law (Eq. 2) that $\alpha$ represents the LLM quality parameter and $\beta$ represents the visual token quality parameter, both denoting the rate at which they influence the downstream error respectively. From Figure 1, we observe that for our selection of language model family (Qwen-1.5) and token compression algorithm, $\alpha=0.077$ is more than five times larger than $\beta=0.015$, signifying that VLM error increases significantly faster when reducing the LLM parameters compared to reducing the number of visual tokens. Therefore, when minimizing inference FLOPs, it is more effective to prioritize reducing visual tokens ($V$) first as the impact on performance is less pronounced than reducing the LLM parameters ($N$).
@@
Figure 2: Our scaling laws (fitted on 0.5-7B VLMs), estimate the performance of 14B VLM with an error margin of less than 2%.
@@
We evaluate the accuracy of our scaling laws (fitted on VLMs of 0.5B-7B range) for predicting the performance for larger models. We estimate the performance of Qwen-1.5 14B using our fitted scaling laws. Our scaling laws estimate the performance with an error margin of less than 2%, as visualized in Figure 2 and Figure 7(b). The log-linear relationship between the error and number of visual tokens persists, and the greater influence of the LLM’s size compared to visual tokens on performance continues to hold. Thus, for VLMs using 7B language model backbones, it is still optimal to increase LLM size to 14B while reducing visual token count for fixed inference costs.
3.3.1 Compute-Optimal Visual Reasoning Inference Favors More LLM Parameters
Observe the pareto optimal curve (black dotted curve) in Figure 1. Our results reveal a striking insight: under a fixed inference budget, visual reasoning tasks heavily favor increasing LLM parameter count. At any given inference FLOPs (denoted by any potential vertical line in Figure 1), the compute-optimal strategy is to allocate FLOPs towards a larger LLM by reducing the number of visual tokens (even to one if the prompt can be pre-computed and cached, $Q=0$).
A similar trend holds in the $Q=50$ regime, where the optimal number of visual tokens is around 16, a 97% reduction from the original number of tokens. This can be intuitively explained by the fact that since the language model is already dedicating a fixed amount of compute to process text input tokens, the initial increase in the number of visual tokens represents only a minor amount of the fraction of compute already being spent. Thus, the optimal tokens increases to 16, compared to just 1 in the cached regime.
In Figure 4, we compare VLMs with varying combinations of LLM size and visual token counts under a fixed inference budget. We observe that for many visual reasoning tasks, increasing the size of the language model while reducing visual tokens can lead to significant relative gains. This may be in part due to the scaling properties of the LLMs themselves, leading to models with stronger world views that can better extrapolate with less visual information than their smaller counterparts (Radford et al., 2021; Wei et al., 2022). We note that this trade-off, while effective for visual reasoning, does not extend to certain tasks, e.g., document comprehension, text identification, etc., where a limited number of tokens may fail to capture the high density of information. We discuss this further in Section 3.4.
Scaling Inference Compute by Simply Repeating Tokens:
Many recent works around scaling test-time compute by introducing special tokens (Goyal et al., 2024a) or multiple parallel generations (Zelikman et al., 2024) have shown promising gains in reasoning tasks for language models. We test this notion with VLMs by repeating the visual input tokens (compressed to 4) multiple times. However, we do not observe any performance gains. This is most likely due to the downstream tasks for VLMs being not as reasoning-intensive, thus highlighting the importance of developing better token compression algorithms and potentially introducing more challenging benchmarks.